OpenStudy (anonymous):

In a parking lot, there are 25 cars, 14 vans, and 11 SUVs. Find the ratio of cars to total vehicles. 25/14 25/11 1/2 2/1

OpenStudy (andriod09):

all you do is you take the 25, then add all the cars, which would be:\[20+14+11=50\] then you would put the 25 on top of the 50, \[\frac{25}{50}\] then you would simplify. \[\frac{25}{50}\div\frac{25}{25}=\frac{1}{2}\]

OpenStudy (anonymous):

thanks:)

OpenStudy (andriod09):

No Problem.